OpenStudy (anonymous):

Find the sum of the series from n=0 to infinity: {[1/(5^n)] + [1/(4^n)]}

OpenStudy (anonymous):

\[ \sum_{n=0}^\infty r^n =\frac 1 {1-r} ,\quad |r| < 1 \]

OpenStudy (anonymous):

your series is the sum of two series like the above with different r.

OpenStudy (anonymous):

yeah I see that now .Thanks!

OpenStudy (swissgirl):

wait can I be annoying?

OpenStudy (swissgirl):

How did u get it into that notation?

OpenStudy (anonymous):

it is the defination

OpenStudy (swissgirl):

ya i see it

OpenStudy (swissgirl):

Thanks